Given that Sandy can produce 10 economics reports or 2 sales calls and Tim can produce 2 economics reports or 1 sales call, which of the following would NOT be a mutually agreeable exchange price for Sandy and Tim?

       
a 1 economics report for 1 sales call
b 1 sales call for 3 economics reports   
c 1 sales call for 4 economics reports
d 1 economics report for 1/4 of a sales call
Answer #1

Sandy can produce 10 economics reports or 2 sales call.

Tim can produce 2 economic reports or 1 sales call.

Opportunity cost of producing 1 economic report in terms of sales call Opportunity cost of producing 1 sale call in terms of economic report
Sandy (2/10) = 0.2 sale call (10 /2) = 5 economic reports
Tim (1/2) = 0.5 sale call (2/1) =2 economic report

The opportunity cost of producing 1 economic report is lower for Sandy than Tim. It means Sandy has comparative advantage in the production of economic report. Hence, Sandy will export economic report

In the similar fashion, Tim has the comparative advantage in the production of sale call. Hence, Tim will export Sale call.

The mutually agreeable price for economic report is a given below.

0.2 sales call < 1 economic report < 0.5 sale call.

Sandy will not export 1 economic report for less than 0.2 sales call and Tim will not be willing to give more than 0.5 sales call in return of 1 economic report.

The mutually agreeable price for sale call is given below:

2 economic report < 1 sale call < 5 economic report.

Tim will not export 1 sale call for less than 2 economic report and Sandy will not be willing to give more than 5 economic report in return of 1 sale call.

Therefore, 1 economic report for 1 sale call would not be mutually agreeable exchange price for Sandy and Tim.

Answer: Option (A)